Ferris Company began January with 4,000 units of its principal product. The cost of each unit is $6. Merchandise transactions for the month of January are as follows: Purchases Date of Purchase Units Unit Cost* Total Cost Jan. 10 3,000 $ 7 $ 21,000 Jan. 18 4,000 8 32,000 Totals 7,000 53,000 * Includes purchase price and cost of freight. Sales Date of Sale Units Jan. 5 2,000 Jan. 12 1,000 Jan. 20 3,000 Total 6,000 5,000 units were on hand at the end of the month. 4. Calculate January's ending inventory and cost of goods sold for the month using Average cost, periodic system.
Answer #1

Calculation of average cost:

Total cost= (4000*6)+53000= 24000+53000= 77000

Total units= 4000+7000= 11000 units

Average cost= total cost/ total number of units= 77000/11000= $7

Ending inventory= 5000*7= $35000

Ending inventory= $35000

Cost of goods sold= 6000*7= $42000

Cost of goods sold= $42000